### v1

# UniUIR: Considering Underwater Image Restoration as An All-in-One Learner

Abstract
Existing underwater image restoration (UIR) methods generally only handle color distortion or jointly address color and haze issues, but they often overlook the more complex degradations that can occur in underwater scenes. To address this limitation, we propose a Universal Underwater Image Restoration method, termed as UniUIR, considering the complex scenario of real-world underwater mixed distortions as an all-in-one manner. To decouple degradation-specific issues and explore the inter-correlations among various degradations in UIR task, we designed the Mamba Mixture-of-Experts module. This module enables each expert to identify distinct types of degradation and collaboratively extract task-specific priors while maintaining global feature representation based on linear complexity. Building upon this foundation, to enhance degradation representation and address the task conflicts that arise when handling multiple types of degradation, we introduce the spatial-frequency prior generator. This module extracts degradation prior information in both spatial and frequency domains, and adaptively selects the most appropriate task-specific prompts based on image content, thereby improving the accuracy of image restoration. Finally, to more effectively address complex, region-dependent distortions in UIR task, we incorporate depth information derived from a large-scale pre-trained depth prediction model, thereby enabling the network to perceive and leverage depth variations across different image regions to handle localized degradation. Extensive experiments demonstrate that UniUIR can produce more attractive results across qualitative and quantitative comparisons, and shows strong generalization than state-of-the-art methods.
Index Terms:
Underwater image restoration, all-in-one manner, Mixture-of-Experts, degradation prior, depth information.
I Introduction
Underwater images have proliferated across fields such as marine biology, underwater exploration, and underwater archaeology[1, 2]. They play a vital role in tasks like water body classification and ocean condition monitoring[3, 4], and in providing immersive visuals of underwater heritage and landscapes[5]. However, due to wavelength-dependent light attenuation and scattering by marine microorganisms[6], underwater scenes often exhibit multiple overlapping distortions, including haze, color shifts, low contrast, and blurred details. In addition, certain adverse imaging conditions, such as low-light environments, can further exacerbate the degradation of underwater image quality. These complex, mixed distortions pose significant challenges for both machine vision tasks and human perception, underscoring the need for effective underwater image restoration (UIR) task.
Figure 1: The figures present a subjective statistical analysis of the predominant distortions in the UIEB [7] dataset. Although each image may exhibit multiple distortions, for simplified classification, each is categorized by its most visually prominent distortion. For example, while Sample 1 shows both color distortion and foreground blurring (blue box) as well as regional blurring in the background (orange box), it is primarily classified as blurred because the blurring is more pronounced than the color distortion, which is less severe compared to other samples like Sample 3. Reference images from the UIEB dataset are provided alongside each sample for comparison.
Due to the absorption of light by water, color distortion is present to varying degrees in every underwater image captured. However, in real-world scenarios, addressing only color distortion is insufficient. As shown in Fig. 1, we organized a group of experienced volunteers to conduct a subjective classification of distortions within the UIEB [7] dataset. We categorized the primary distortions in each underwater scene image into five major types: color cast, low contrast, haze, low light, and blur. It is important to note that when an image is primarily classified under one type of distortion, such as blur, it does not mean that only this type of distortion is present. For instance, as indicated by the orange and blue boxes in Sample 1 of Fig. 1, an image predominantly affected by blur may also exhibit color distortion in foreground objects. This highlights the extensive variety of distortion types found in underwater scenes, where multiple distortions often co-occur. Given that the underwater environment poses increasing challenges, this observation underscores the urgent need to develop comprehensive underwater restoration methods capable of handling multiple types of distortions simultaneously.
Figure 2: Comparison between our proposed UniUIR and previous deep learning-based approaches to underwater image restoration. (a) Previous depth estimation methods rely on costly paired ground-truth depth data, while physical model-based approaches fail in complex distortions due to idealized assumptions, leading to poor generalization. (b) Our method leverages a depth prediction network trained on large-scale datasets and a reliable degradation prior extractor, demonstrating improved generalization in complex distortion scenarios through a mixture-of-experts (MoE)-based network.
Existing UIR methods can be broadly classified into three categories: visual prior-based, physical model-based, and deep learning-based approaches. Visual prior-based methods focus on adjusting image pixel values such as contrast, brightness, and saturation [8, 9], but fail to account for the physical degradation processes specific to underwater environments. Physical model-based methods [9, 10, 11] mainly estimate physical medium parameters of the underwater imaging process, yet their adaptability is limited by environmental complexity, resulting in reduced robustness. Recently, deep learning-based approaches [7, 12, 13, 14, 15, 16] have emerged, demonstrating notable improvements through end-to-end optimization. However, these methods often assume that all distortions can be handled simultaneously, typically focusing on specific types of distortions while neglecting the interactions and conflicts between them. As a result, their performance is limited in complex scenes with multiple simultaneous distortions. As illustrated in Fig. 2(a), many of these methods rely on physical modeling [17, 18] or jointly optimized depth estimation networks [14, 19, 20], which present several challenges: 1) Depth estimation networks require paired ground-truth depth data for optimization, which is difficult and costly to obtain. 2) Physical models rely on idealized assumptions, making them potentially inadequate for handling the complexity of real-world underwater environments, often fail to handle distortions such as detail blurring or low-light conditions. In contrast, our method (Fig. 2(b)) emphasizes representing degradation priors and leverages the capabilities of large-scale pre-trained depth prediction models. By incorporating an encoder-decoder framework equipped with a mixture-of-experts, our approach effectively addresses multiple distortions, including extreme low-light underwater scenarios, demonstrating robust and comprehensive restoration performance.
In this study, we aim to systematically address the aforementioned challenges. Specifically, we propose a novel universal underwater image restoration method, termed UniUIR, designed to tackle the complex real-world scenario of mixed underwater distortions in an all-in-one manner. By analyzing the distortion characteristics and distributions in existing underwater image datasets, we develop MMoE-UIR, a Mamba Mixture-of-Experts-based restoration backbone that effectively captures and addresses diverse degradation patterns. To further enhance restoration performance, we design a spatial-frequency prior generator (SFPG) capable of extracting degradation priors from both spatial and frequency domains while adaptively selecting the most relevant task-specific prompts based on image content. Additionally, to address inconsistencies between distortions across different image regions, we incorporate scene depth information using the Depth Anything V2 [21] model, enabling the network to better distinguish and focus on visual discrepancies between foreground and background elements. Finally, to recover fine-grained details and mitigate information loss inherent in degradation priors, we leverage the powerful generative capabilities of diffusion models [22, 23]. The latent condition diffusion model is introduced to refine the restoration process, effectively modeling complex distributions and enabling the reconstruction of high-quality images with accurate and visually consistent details.
In summary, our main contributions are as follows.
❑ We are among the first to address the complex scenario of underwater mixed distortions from an integrated perspective. To tackle this challenge, we propose UniUIR, a novel all-in-one method for underwater image restoration, designed to achieve more authentic restoration results.
❑ To leverage the interdependencies among various degradations in underwater environments, we developed the Mamba Mixture-of-Experts module. This module facilitates each expert in recognizing specific degradations and collectively inferring task-specific priors, while preserving both local and global feature representations.
❑ We leverage the large-scale pre-trained depth prediction model to incorporate scene depth information, enabling the identification of distortion inconsistencies across different image regions. Additionally, we introduce the spatial-frequency prior generator, which adaptively selects the most appropriate task-specific prompts to extract degradation priors.
❑ Compared to state-of-the-art underwater image restoration methods, UniUIR achieves superior results in extensive qualitative and quantitative evaluations. Furthermore, UniUIR demonstrates strong generalization capabilities across other visual tasks.
II Related Work
II-A Underwater Image Restoration
Existing underwater image restoration (UIR) methods can be broadly divided into traditional and deep learning-based approaches. Traditional methods are further categorized into non-physical model-based and physical model-based techniques. Non-physical model-based methods focus on enhancing image quality by adjusting pixel properties such as contrast and saturation [24, 25, 8], without considering the optical transmission process. While computationally efficient and straightforward, these methods often lack a physical understanding of underwater light propagation, leading to issues such as over- or under-saturation in complex scenes. Physical model-based methods, on the other hand, rely on mathematical formulations of the degradation process, accounting for factors like light scattering and medium transmission. While effective in controlled settings, their reliance on predefined models often limits their adaptability to diverse underwater environments, reducing robustness and generalizability.
Recently, deep learning–based methods [7, 12, 13, 15, 26, 18, 19, 16, 27, 28] have shown significant potential in underwater image restoration (UIR). These methods leverage advanced neural network architectures and techniques to address challenges in image degradation. For instance, Waternet [7] integrates enhanced input modes with confidence maps to generate improved images, while Ucolor [12] employs multi-color space embedding guided by medium transmission to enhance detail and color restoration. Semi-UIR [26] combines contrastive and semi-supervised learning to improve performance and robustness, and HCLR-net [16] uses hybrid contrastive learning with locally randomized perturbations to enhance image quality. Despite these advancements, deep learning-based methods often assume that all distortions can be addressed simultaneously. This assumption leads to a focus on main distortion, for example, color cast or low contrast, while overlooking interactions and conflicts between multiple distortions. Consequently, their performance remains limited in complex underwater scenes with diverse and simultaneous degradations.
II-B All-in-One Image Restoration
All-in-one image restoration [29, 30, 31, 32, 33, 34] has emerged as a promising approach in low-level vision tasks, aiming to recover clean images from multiple degraded inputs using a unified model. Compared to task-specific [35, 36, 37, 38, 39, 40] and general [41, 42, 43, 44, 45] restoration methods, all-in-one restoration offers advantages in model storage efficiency and practical applications. The main challenge lies in designing a single architecture capable of addressing diverse degradation types. To tackle this, methods like AirNet [29] employ contrastive learning for discriminative degradation representations, while IDR [30] uses a two-stage, ingredients-oriented approach. Vision prompts, as used in PromptIR [31] and ProRes [32], further enhance the handling of diverse degradations. More recently, MPerceiver [34] has leveraged pre-trained large-scale vision models to excel in all-in-one restoration tasks.
Underwater Image Restoration (UIR) naturally aligns with the all-in-one restoration framework due to the presence of multiple distortions in underwater scenes (as shown in Fig. 1). However, directly applying general all-in-one methods to underwater images is suboptimal because of the significant domain shift between natural and underwater environments. To address this gap, we propose a specialized all-in-one restoration approach tailored for UIR. By introducing a mixture-of-experts mechanism, our method adopts a “divide-and-conquer” strategy, dynamically allocating expert submodules to handle specific types of distortions, thereby improving restoration precision and adaptability. Additionally, our method integrates physical depth information and emphasizes spatial and frequency domain prior representations, enabling it to better capture the unique characteristics of underwater imagery. These innovations allow our approach to achieve superior restoration results, particularly in challenging underwater environments with complex and overlapping distortions.
II-C Diffusion Models for Image Restoration
Diffusion models (DM) [22, 23], initially developed for generative tasks, have demonstrated strong potential in image restoration by addressing inverse problems through progressive denoising. Their capability to model complex data distributions has led to state-of-the-art results in low-level vision tasks such as super-resolution [46, 47] and low-light image enhancement [48, 49]. Recently, the generative power of DM has attracted attention in underwater image restoration (UIR) [50, 51]. For example, Tang et al. [50] proposed a diffusion-based underwater image enhancement method, while UIEDP [51] introduced a framework incorporating diffusion priors, formulating UIR as posterior sampling conditioned on degraded inputs.
Although DM excel at generating high-quality images, they face notable challenges in UIR, including low efficiency, limited adaptability to diverse degradation types, insufficient integration of environmental priors, and constrained generalization capabilities. To address these limitations, our UniUIR incorporates a latent condition diffusion model to refine the restoration process and reconstruct high-quality images with consistent and precise details, offering a more effective solution for UIR.
Figure 3: Overview of the proposed UniUIR. (a) In Stage I, the depth map $\mathbf{D}_{lq}$, degradation prior $\mathbf{Z}$, and image $\mathbf{X}_{lq}$ are processed by the MMoE-UIR network to reconstruct the high-quality image $\mathbf{X}_{hq}$. The $\mathcal{L}_{1}$ loss and the edge-aware depth loss $\mathcal{L}_{\text{depth}}$ encourage the network to focus on overall pixel-level reconstruction and the accurate recovery of edges and structures, respectively. (b) In Stage II, the prior $\mathbf{Z}$ extracted by the SFPG and conditional embedding $\mathbf{C}$ will undergoes Latent Condition Diffusion Model to iteratively denoise, yielding $\mathbf{\hat{Z}}$. Then, $\mathbf{D}_{lq}$, $\mathbf{X}_{lq}$, and $\mathbf{\hat{Z}}$ are processed by MMoE-UIR to restore the clean image. The network is fine-tuned with $\mathcal{L}_{1}$ and $\mathcal{L}_{\text{Diff}}$ to optimize image restoration and the diffusion process. (c) During inference, SFPG* first produces the conditional embedding $\mathbf{C}$. Subsequently, a random noise $\mathbf{\hat{Z}_{t}}$ is sampled from a Gaussian distribution. Through the reverse process, $\mathbf{C}$ and $\mathbf{\hat{Z}_{t}}$undergo iterative denoising to generate $\mathbf{\hat{Z}}$. Lastly, the fine-tuned MMoE-UIR then leverages $\mathbf{X}_{lq}$, $\hat{\mathbf{Z}}$, and $\mathbf{D}_{lq}$ to reconstruct the final high-quality image $\mathbf{X}_{hq}$.
III Method
III-A Overview Pipeline
The proposed UniUIR framework comprises three stages: training stage I, training stage II, and inference stage. As illustrated in Fig. 3, during training stage I, given an underwater image $\mathbf{X}_{lq}\in\mathbb{R}^{H\times W\times 3}$, a pre-trained Depth Anything V2 model is used to incorporate scene depth information $\mathbf{D}_{lq}$ from $\mathbf{X}_{lq}$. Meanwhile, the reference image $\mathbf{X}_{gt}$ is concatenated with $\mathbf{X}_{lq}$ and passed through the spatial-frequency prior generator (SFPG) to obtain a degradation prior $\mathbf{Z}\in\mathbb{R}^{\hat{C}}$. Subsequently, $\mathbf{D}_{lq}$ and $\mathbf{Z}$ together with $\mathbf{X}_{lq}$ are processed by the MMoE-UIR, an $n$-stage U-shaped network, to reconstruct high-quality result $\mathbf{X}_{hq}$. In training stage II, the prior $\mathbf{Z}$ extracted by the pre-trained SFPG undergoes a forward diffusion process to generate the noised $\mathbf{Z}_{t}$, and the $\mathbf{X}_{lq}$ is passed through SFPG* to yield a conditional embedding $\mathbf{C}$. During the reverse process, $\mathbf{Z}_{t}$ and $\mathbf{C}$ are fed into a UNet–based denoising network to iteratively perform denoising operations, yielding $\mathbf{\hat{Z}}$. Finally, $\mathbf{D}_{lq}$, $\mathbf{X}_{lq}$, and $\mathbf{\hat{Z}}$ are processed by MMoE-UIR to restore a clean underwater image. Throughout this process, MMoE-UIR is fine-tuned based on the initial training stage. The optimization employs both pixel loss to constrain the overall image restoration and Diff loss to regulate the diffusion model’s generation process. During inference stage, given $\mathbf{X}_{lq}$, SFPG* first produces the conditional embedding $\mathbf{C}$. Subsequently, a random noise $\mathbf{\hat{Z}_{t}}$ is sampled from a Gaussian distribution. Through the reverse process, $\mathbf{C}$ and $\mathbf{\hat{Z}_{t}}$undergo iterative denoising to generate $\mathbf{\hat{Z}}$. Lastly, the fine-tuned MMoE-UIR then leverages $\mathbf{X}_{lq}$, $\hat{\mathbf{Z}}$, and the extracted depth $\mathbf{D}_{lq}$ to reconstruct the final high-quality image $\mathbf{X}_{hq}$. In the following subsections, we will provide a detailed description of the roles and processes of these modules.
Optimization. In the first stage, we use $\mathcal{L}_{1}$ loss as the pixel-level loss to constrain the overall learning process, in addition, we introduce edge-aware depth loss $\mathcal{L}_{\text{depth}}$ to let the network focus on the gradient information of the depth map, emphasizing the accurate recovery of edges and structures. The training loss can be describe as follows:
$\mathcal{L}_{\text{stageI}}=\mathcal{L}_{{1}}(\mathbf{X}_{gt},\mathbf{X}_{hq})+\lambda_{1}\mathcal{L}_{\text{depth}},$ (1)
$\mathcal{L}_{\text{depth}}=\mathcal{L}_{1}(\mathbf{D}_{pseudo},\mathbf{D}_{hq})+\lambda_{2}\mathcal{L}_{\text{grad}}(\mathbf{D}_{pseudo},\mathbf{D}_{hq}),$ (2)
$\mathcal{L}_{\text{grad}}=\left(\left|\frac{\partial({\mathbf{D}_{{pseudo}}-\mathbf{D}_{hq}})}{\partial x}\right|+\left|\frac{\partial({\mathbf{D}_{{pseudo}}-\mathbf{D}_{hq}})}{\partial y}\right|\right),$ (3)
where $\mathbf{X}_{gt}$ and $\mathbf{X}_{hq}$ denote the ground-truth image and the restored high-quality image, respectively. Similarly, $\mathbf{D}_{pseudo}$ and $\mathbf{{D}}_{hq}$ represent the pseudo depth and high-quality depth maps derived from the Depth Anything V2 Model using $\mathbf{X}_{gt}$ and $\mathbf{X}_{hq}$, respectively. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ are introduced to balance the contributions of individual terms in the total loss function. Our training strategy incorporates both edge-aware depth loss and $\mathcal{L}_{1}$ loss to achieve complementary benefits. The edge-aware depth loss enables the model to capture precise edge information from the pseudo ground-truth depth, while the $\mathcal{L}_{1}$ loss ensures accurate overall depth estimation. This dual-loss approach enhances depth prediction performance by effectively combining edge precision with global consistency.
In the second stage, we jointly optimize using the diffusion loss $\mathcal{L}_{\text{diff}}$ and pixel-level loss. The training loss can be described as follows:
$\mathcal{L}_{\text{stageII}}=\mathcal{L}_{{1}}(\mathbf{X}_{gt},\mathbf{X}_{hq})+\mathcal{L}_{\text{diff}},\enspace\mathcal{L}_{\text{diff}}=\mathcal{L}_{{1}}(\mathbf{Z},\mathbf{\hat{Z}}).$ (4)
III-B MMoE-UIR
As depicted in Fig. 4, the low-quality underwater input $\mathbf{X}_{lq}$ is processed by a $3\times 3$ convolution-based feature extraction module, followed by a four-stage U-shaped encoder–decoder network. The core component, the Mamba Mixture of Experts Block (MMoEB), integrates the Vision State-Space Module (VSSM) and Water Mixture-of-Experts (W-MoE). Each MMoEM contains multiple MMoEBs. By leveraging task-specific degradation priors and depth information, the MMoEB refines these priors to facilitate multi-scale feature fusion and reconstruction, improving adaptability to complex underwater environments and overall restoration performance.
Mamba Mixture-of-Experts Block. As shown in Fig. 4(b), given the deep feature $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$, we first use the Layer Norm (LN) followed. Similarly, we allow the prior feature $\mathbf{z}$ to interact with the normalized feature before feeding it into the VSSM. This process can be represented as:
$\mathbf{F}_{d}=\text{VSSM}\big{(}\text{LN}(\mathbf{F})\odot\text{Linear}(\mathbf{Z}_{n})+\text{Linear}(\mathbf{Z}_{n})\big{)}+\mathbf{F},$ (5)
where $\mathbf{Z}_{n}$ represents the $n$-th stage compression feature, which is obtained by downsampling the original feature $\mathbf{Z}$ for $n$ times. The generated deep features $\mathbf{F}_{d}$ are subsequently fed into the FFN model, where they are first enhanced through LN. The normalized features then interact with the enhanced deep features, after which they are passed into the W-MoE. This process can be expressed as:
$\mathbf{\hat{F}}=\mathcal{M}\Big{(}\text{LN}(\mathbf{F}_{d})\odot\text{Softmax}\big{(}\mathcal{C}_{3\times 3}(\mathbf{D}_{lq}^{n})\big{)}\Big{)}+\mathbf{F}_{d},$ (6)
where $\mathbf{\hat{F}}$ represents the output of the MMoEB; $\mathcal{M}(\cdot)$ denotes the Water Mixture-of-Experts operation; $\mathcal{C}_{3\times 3}$ denotes 3$\times$3 convolution, followed by a ReLU function; $\mathbf{D}_{lq}$ is generated from the pre-trained Depth Anything V2 model. Each MMoEM contains multiple MMoEBs.
Figure 4: The detailed structure of the proposed MMoE-UIR. (a) Mamba Mixture-of-Experts Underwater Image Restoration (MMoE-UIR); (b) Mamba Mixture-of-Experts Block (MMoEB); (c) Water Mixture-of-Experts (W-MoE).
Water Mixture-of-Experts. As illustrated in Fig. 4(c), given an input feature $\mathbf{F}_{m}\in\mathbb{R}^{H\times W\times C}$ to W-MoE, we employ a $3\times 3$ convolution for feature projection and subsequently split the features along the channel dimension to create two distinct views $\mathbf{F}_{m}^{a}$ and $\mathbf{F}_{m}^{b}\in\mathbb{R}^{H\times W\times C}$. Simultaneously, a parallel depth-wise convolution extracts the local spatial context $\mathbf{\hat{F}}_{m}^{a}$ before feeding both extracted feature maps into the Mixture-of-Experts. To further investigate the intricacies of inter-dependencies among the extracted features while reducing complexity, we implement low-rank decomposition for the inputs while modeling global contextual relationships. A single low-rank expert $\mathcal{E}_{i}$ can be formulated as:
$\mathcal{E}_{i}=\mathcal{T}_{C_{l}\rightarrow C_{h}}^{3}\left(\mathcal{T}_{C_{h}\rightarrow C_{l}}^{1}(\mathbf{{F}}_{m}^{a})\odot\mathcal{T}_{C_{h}\rightarrow C_{l}}^{2}(\mathbf{\hat{F}}_{m}^{b})\right),$ (7)
here, for the $i$-th expert $\mathcal{E}_{i}$, the number of channels $C_{h},C_{l},C$ are related as follows: $C_{h}=C$ and $C_{l}=2^{i+1}$. The linear transfer layer, denoted as $\mathcal{T}(\cdot)$, is implemented using $1\times 1$ point-wise convolutions that compress the encoded features along the channel dimension to their low-rank approximation $R_{i}$, where $i\in{1,\cdots,N}$. After modulating spatial cues through element-wise multiplication with contextual cues in the low-dimensional space, another linear transfer layer, $\mathcal{T}_{C_{l}\rightarrow C_{h}}^{3}$, restores the features back to their original dimension $C$. This reconstruction enables the effective extraction of channel-wise spatial content while implicitly integrating critical spatial and channel dependencies. Additionally, we employ a routing network $\mathcal{G}$, composed of two MLP layers. This network systematically explores the search space to identify the optimal low-rank expert based on the input features and the network depth. The final output, $\mathbf{\hat{F}}_{m}$, of the W-MoE is formulated as follows:
$\mathbf{\hat{F}}_{m}=\sum_{i=1}^{N}\mathcal{G}(\mathbf{{\hat{F}}}_{m}^{b})\cdot\mathcal{E}_{i}(\mathbf{{F}}_{m}^{a},\mathbf{{\hat{F}}}_{m}^{b}),$ (8)
where $\mathcal{G}(\cdot)$ represents the learned router module, while $\mathcal{E}_{i}(\cdot)$ denotes the output of the $i$-th expert. The sparsity inherent in the router function $\mathcal{G}(\cdot)$ optimizes computational efficiency by assigning higher weights to the top-$k$ low-rank experts. During training, the model learns from all experts; during inference, only the selected top-$k$ experts are utilized for computation, significantly enhancing efficiency. For clarity, the pseudocode of the W-MoE is provided in Algorithm 1.
III-C Spatial-Frequency Prior Generator
Reconstructing image details in severely degraded regions is challenging due to the significant loss of high-frequency information caused by complex distortions. To restore these missing details, external guidance from high-quality underwater images is necessary. To achieve this, we use a Spatial-Frequency Prior Generator (SFPG) to extract a compact prior $\mathbf{Z}$ from $\mathbf{X}_{lq}$ and $\mathbf{X}_{gt}$, injecting it into MMoE-UIR to provide correct information for restoring the lost details. Specifically, as shown in Fig. 3, SFPG first employs PixelUnshuffle and 3$\times$3 convolutions with LReLU activation functions to downsample the concatenated images $\mathbf{X}_{c}\in\mathbb{R}^{H\times W\times 6}$ and capture effective information, resulting in $\mathbf{X}_{s}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C^{\prime}}$. Subsequently, two parallel branches of three residual blocks with 3$\times$3 convolutions and LReLU activation functions are utilized to extract features from the amplitude and phase of $\mathbf{X}_{s}$.
$\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}=\mathcal{F}(\mathbf{X}_{s}),\enspace\tilde{\mathbf{X}}_{\mathcal{A}},\tilde{\mathbf{X}}_{\mathcal{P}}=\text{ResBlock}(\mathbf{X}_{\mathcal{A}},\mathbf{X}_{\mathcal{P}}),$ (9)
$\mathbf{X}_{f}=\mathcal{F}^{-1}(\tilde{\mathbf{X}}_{\mathcal{A}},\tilde{\mathbf{X}}_{\mathcal{P}}),$ (10)
Where $\mathcal{F}(\cdot)$ denotes the 2D FFT and $\mathcal{F}^{-1}(\cdot)$ represents the 2D IFFT, ${\mathbf{X}}_{\mathcal{A}}$ and ${\mathbf{X}}_{\mathcal{P}}$ respectively denote the amplitude and phase features. After being mapped back to the spatial domain, these features are refined by three convolutional layers with LReLU activation functions and a global average pooling (GAP), producing a compact representation.
To effectively accommodate various degradation types of underwater image restoration tasks within a unified model, we embed Task-Related Prompts (TRP) at the end of the SFPG, generating prompt information closely tied to specific tasks. Concretely, we employ the softmax output $\mathcal{S}\in\mathbb{R}^{N}$ as the weighting factor for the basic prompts associated with each degradation, to combine and construct prompts for that particular task. Let $\mathcal{P}\in\mathbb{R}^{N\times\hat{C}}$ be the basic task-related prompts. The final generated prior $\mathbf{Z}$ can be expressed as:
$\mathbf{Z}={\mathcal{P}}\otimes\mathcal{S},$ (11)
where $\otimes$ represents matrix multiplication.
Algorithm 1 Water Mixture-of-Experts
1:  Input: Input feature $\mathbf{{F}}_{m}$, separated feature $\mathbf{{F}}_{m}^{a}$, extracted local features $\mathbf{{\hat{F}}}_{m}^{b}$
2:  Parameters: $N$ Experts $\mathcal{E}$, Router $\mathcal{G}$, and top-k expert
3:  Compute router outputs: $\mathbf{g}=\mathcal{G}(\mathbf{\hat{x}}_{a})$
4:  Normalize weights: $\mathbf{w}=\text{Softmax}(\mathbf{g})$
5:  Select top-k expert: $w_{1},w_{2},...,w_{k}=\text{top-k}(\mathbf{w},k)$
6:  if training then
7:     for each $\mathcal{E}_{i}\in\mathcal{E}$ do
8:        $\mathbf{y}^{i}_{\mathcal{E}_{i}}=\mathcal{T}_{C_{l}\rightarrow C_{h}}^{3}\left(\mathcal{T}_{C_{h}\rightarrow C_{l}}^{1}(\mathbf{{F}}_{m}^{a})\odot\mathcal{T}_{C_{h}\rightarrow C_{l}}^{2}(\mathbf{\hat{F}}_{m}^{b})\right)$
9:     end for
10:     Compute final output: $\mathbf{y}=\sum_{i=1}^{N}w_{i}\cdot\mathbf{y}^{i}_{\mathcal{E}_{i}}$
11:  else
12:     Compute final output: $\mathbf{y}=w_{\text{top-k}}\cdot\mathbf{y}^{\text{top-k}}_{\mathcal{E}_{i}}$
13:  end if
14:  Output: Final output $\mathbf{y}$
III-D Latent Conditional Diffusion Model
Forward Diffusion Process. As illustrated in Fig. 4(b), we first utilize the SFPG trained in the first stage to generate $\mathbf{Z}$. Subsequently, a diffusion process is applied to gradually add Gaussian noise to $\mathbf{Z}$ over $T$ time steps, resulting in the noisy representation $\mathbf{Z}_{T}$, which can be expressed as:
$q(\mathbf{Z}_{T}\mid\mathbf{Z})=\mathcal{N}(\mathbf{Z}_{T};\sqrt{\bar{\alpha}_{T}}\mathbf{Z},(1-\bar{\alpha}_{T})\mathbf{I}),$ (12)
where $T$ is total number of iterations; $\alpha_{T}=1-\beta_{T}$ and $\bar{\alpha}_{T}=\prod_{i=1}^{T}\alpha_{i}$. ${\beta}_{t}(t=1,\cdots,T)$ is hyper parameter controlling the variance of the noise; $\mathcal{N}$ denotes the Gaussian distribution.
Reverse Diffusion Process. The reverse process runs backwards from $\mathbf{Z}_{T}$ to $\mathbf{\hat{Z}}$. For the reverse step from $\mathbf{Z}_{t}$ to $\mathbf{Z}_{t-1}$, we use the posterior distribution as:
$p(\mathbf{Z}_{t-1}\mid\mathbf{Z}_{t},\mathbf{\hat{Z}})=\mathcal{N}(\mathbf{Z}_{t-1};\bm{\mu}_{t}(\mathbf{Z}_{t},\mathbf{\hat{Z}}),\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}\mathbf{I}),$ (13)
$\bm{\mu}_{t}(\mathbf{Z}_{t},\mathbf{\hat{Z}})=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{Z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}),$ (14)
[…]
$\mathbf{Z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{Z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{Z}_{t},\mathbf{C},t)\right)+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t},$ (15)
where ${\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})$. By iteratively sampling $\mathbf{Z}_{t}$ using Eq. (15) $T$ times, we can generate the predicted prior feature $\mathbf{\hat{Z}}\in\mathbb{R}^{N\times\hat{C}}$. The predicted prior feature is then used to guide MMoEB in Fig. 4(b).
TABLE I: Datasets summary for underwater image restoration task.
Setting Train/Test Dataset Type No. of Images Main Objects Resolution
Set Training UIEB-T[7] Paired 800 Wreck, People, Fish, Reef, and Coral (259$\times$194), $\cdots$, (2180$\times$1447)
[…]
U45[52] Unpaired 45 Wreck, People, Fish, Reef, and Coral 256$\times$256
SQUID-16[53] Unpaired 16 Wreck and Reef 512$\times$512
Challenge-60[7] Unpaired 60 People, Fish, Reef, Medusa (259$\times$194), $\cdots$, (2000$\times$1124)
UCCS[54] Unpaired 300 Reef and Sea Urchin 400$\times$300
EUVP-330[55] Unpaired 330 People, Fish, Reef, and Coral (320$\times$240), $\cdots$, (960$\times$540)
Set Training LSUI-T[13] Paired 3879 Fish, Reef, Coral, and Sea Urchin (256$\times$256), $\cdots$, (1280$\times$1024)
[…]
IV Experiments
IV-A Experimental Setup
Datasets. As shown in Tab. I, the datasets used in this study include both paired and unpaired sets. For paired training data, we utilize UIEB-T [7] and LSUI-T [13]. Unpaired test datasets include U45 [52], SQUID-16 [53], Challenge-60 [7], UCCS [54], and EUVP-330 [55]. The UIEB dataset consists of 950 real underwater images, 890 of which have corresponding reference images. We use 800 of these paired images for training (UIEB-T) and the remaining 90 for testing (T90). The remaining 60 images, lacking satisfactory reference images, constitute the Challenge-60 subset. The LSUI dataset is a large-scale underwater image dataset comprising 4279 image pairs. Following [13], we use 3879 images for training (LSUI-T) and the remaining 400 for testing (LSUI-400). The U45 dataset presents 45 real-world underwater images with challenges including color casts, low contrast, and haze-like effects. The SQUID dataset contains 57 stereo underwater image pairs, each including a color chart for color restoration evaluation. Following [12], we selected 16 representative samples from SQUID, referred to as SQUID-16. Further details about the datasets are provided in Tab. I.
Evaluation Metrics. We evaluate performance using reference metrics including Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) [56], as well as non-reference metrics such as the Underwater Colour Image Quality Evaluation Metric (UCIQE) [57], Underwater Image Quality Measure (UIQM) [58], and URanker [15]. For the metrics PSNR, SSIM, UCIQE, UIQM, and URanker, higher scores indicate better performance. Conversely, for the LPIPS metric, lower scores are preferable.
Baselines. We compare our method against existing underwater image restoration methods, encompassing both traditional and deep learning approaches. These include a traditional method MLLE [25] and the following deep learning models: WaterNet [7], FUnIE [59], S-UWNet [60], Ucolor [12], PUIE-Net [61], Ushape [13], PUGAN [14], NU2Net [15], Semi-UIR [26], GUPDM[18], HCLR-Net [16]. Additionally, we compared our approach with the all-in-one image restoration method PromptIR[31] and the MambaIR[43] based on the Mamba architecture [62].
Implementation Details. All experiments were conducted using PyTorch on 8 NVIDIA RTX 3090 GPUs and optimized using the AdamW optimizer with momentum parameters (0.9, 0.999). During training stage I and II, UniUIR was trained for 50K and 200K iterations, respectively, with an initial learning rate of $2\times 10^{-4}$, gradually reduced to $1\times 10^{-6}$ using cosine annealing. The batch size was set to 8. MMoE-UIR utilized a 4-level encoder-decoder architecture with [3, 5, 6, 6] MMoEBs and channel dimensions of [32, 64, 128, 256] for levels 1 through 4. The channel numbers ${\hat{C}}$ in both SFPG and SFPG* were set as 256. A linear noise schedule was employed for diffusion process with parameters $\alpha_{1}=0.99$, $\alpha_{T}=0.1$, and a time step of $T=4$. Depth information was extracted using the large version of the Depth Anything V2 model [21]. For fair comparisons, all deep learning-based UIR methods were retrained using the original settings specified in their respective papers. Input images were cropped to 128$\times$128 pixels, and random flipping was applied for data augmentation. The hyperparameters were set as $\lambda_{1}=0.1$ and $\lambda_{2}=0.5$.
Figure 5: Visual comparison of restored results for the T90[7] test set.
TABLE II: Quantitative results on referenced-based datasets: T90 and LSUI-400. The top three results are marked with red, blue, and green, respectively.
Method Source T90 [7] LSUI [13]
[…]
WaterNet [7] TIP’19 20.87 0.911 0.157 0.595 2.823 1.371 23.07 0.853 0.221 0.584 2.807 1.398
FUnIE [59] IRAL’20 19.18 0.865 0.179 0.599 3.002 1.157 22.05 0.832 0.239 0.567 2.681 1.083
S-UWnet [60] AAAI’21 18.37 0.753 0.192 0.581 2.432 1.298 21.33 0.884 0.243 0.591 2.719 1.385
Ucolor [12] TIP’21 20.81 0.904 0.160 0.573 3.019 1.416 23.28 0.890 0.211 0.589 2.804 1.201
MLLE [25] TIP’22 18.68 0.855 0.286 0.598 2.854 1.007 22.23 0.835 0.254 0.569 3.071 1.116
PUIE-Net [61] ECCV’22 22.03 0.891 0.241 0.587 2.946 1.462 23.93 0.904 0.206 0.614 2.981 1.418
Ushape [13] TIP’23 20.97 0.864 0.203 0.579 3.067 1.358 24.46 0.901 0.190 0.599 2.994 1.411
PUGAN [14] TIP’23 22.79 0.922 0.193 0.592 3.104 1.347 25.11 0.915 0.184 0.615 3.112 1.309
NU2Net [15] AAAI’23 22.91 0.922 0.174 0.605 3.153 1.517 25.32 0.910 0.162 0.611 3.005 1.413
Semi-UIR [26] CVPR’23 24.19 0.923 0.151 0.568 2.537 1.479 27.08 0.907 0.134 0.614 3.016 1.435
GUPDM [18] ACM MM’23 24.33 0.928 0.133 0.552 3.012 1.306 27.77 0.917 0.138 0.627 3.103 1.466
PromptIR [31] NeurIPS’23 23.46 0.919 0.152 0.571 2.933 1.376 26.69 0.905 0.141 0.594 2.984 1.321
HCLR-Net [16] IJCV’24 23.72 0.917 0.158 0.575 2.537 1.454 26.98 0.904 0.130 0.597 3.272 1.483
MambaIR [43] ECCV’24 23.57 0.922 0.147 0.584 2.851 1.385 26.32 0.905 0.143 0.602 3.056 1.322
UniUIR (Ours) 25.11 0.933 0.112 0.601 3.117 1.534 28.42 0.926 0.123 0.610 3.138 1.507
TABLE III: Quantitative results on non-referenced-based datasets: U45, SQUID-16, Challenge-60, UCCS, and EUVP-330. The top three results are marked with red, blue, and green, respectively.
Method Source U45 [52] SQUID-16 [53] Challenge-60 [7] UCCS [54] EUVP-330 [55]
[…]
WaterNet [7] TIP’19 0.572 3.195 1.312 0.541 2.318 1.746 0.566 2.653 1.223 0.545 3.058 1.281 0.524 3.019 1.685
FUnIE [59] IRAL’20 0.589 3.181 1.763 0.532 2.417 1.456 0.570 3.158 1.304 0.541 2.986 1.043 0.548 2.942 1.728
S-UWnet [60] AAAI’21 0.533 2.933 1.681 0.491 1.894 1.279 0.466 2.396 1.473 0.486 2.759 1.312 0.523 3.018 1.609
Ucolor [12] TIP’21 0.564 3.151 1.485 0.514 2.215 1.613 0.532 2.746 1.213 0.550 3.019 1.064 0.561 3.114 1.546
MLLE [25] TIP’22 0.593 2.599 1.349 0.531 2.314 1.378 0.581 2.310 1.257 0.544 2.985 1.145 0.568 3.026 1.629
PUIE-Net [61] ECCV’22 0.578 2.799 1.632 0.522 1.923 2.014 0.558 2.521 1.307 0.536 3.003 1.352 0.548 2.979 1.415
Ushape [13] TIP’23 0.553 3.048 1.737 0.528 2.256 1.857 0.534 2.783 1.363 0.567 3.012 1.374 0.557 3.013 1.679
PUGAN [14] TIP’23 0.590 3.095 1.815 0.566 2.399 1.956 0.612 3.001 1.641 0.536 2.977 1.549 0.541 2.858 1.747
NU2Net [15] AAAI’23 0.595 3.206 1.804 0.534 2.480 2.156 0.564 2.907 1.557 0.601 2.994 1.624 0.571 3.127 1.801
Semi-UIR [26] CVPR’23 0.601 3.323 1.831 0.558 2.537 1.945 0.573 2.925 1.685 0.552 3.039 1.321 0.526 2.994 1.864
GUPDM [18] ACM MM’23 0.566 3.057 1.872 0.547 2.459 2.013 0.548 2.687 1.751 0.586 3.157 1.583 0.563 3.014 1.733
PromptIR [31] NeurIPS’23 0.573 2.847 1.736 0.531 2.473 1.834 0.577 2.808 1.638 0.549 2.897 1.512 0.545 2.911 1.844
HCLR-Net [16] IJCV’24 0.585 3.013 1.927 0.538 2.508 2.109 0.569 2.737 1.615 0.541 3.009 1.784 0.559 3.102 2.052
MambaIR [43] ECCV’24 0.589 3.127 1.825 0.553 2.322 1.895 0.552 2.983 1.493 0.564 3.114 1.417 0.537 2.975 1.806
UniUIR (Ours) 0.609 3.278 2.028 0.570 2.514 2.217 0.593 3.218 1.825 0.576 3.215 1.603 0.581 3.198 1.927
Figure 6: Visual comparison of restoration results on the non-reference datasets: U45 [52], SQUID-16 [53], Challenge-60 [7], UCCS [54], and EUVP-330 [55].
Figure 7: Visual results comparison of various network settings. Index A serves as the baseline. Index B extends Index A by incorporating Depth Anything V2. Index C builds upon Index B by integrating the Mamba Mixture-of-Experts Block (MMoEB). Index D further enhances Index C with the Spatial-Frequency Prior Generator (SFPG). Finally, Index E, denoted as UniUIR, represents the full proposed model, augmenting Index D with a Latent Diffusion Model (LCDM).
IV-B Qualitative Comparison
As depicted in Tab. II, some early deep learning-based underwater image restoration (UIR) methods such as FUnIE[59] and S-UWnet[60], which employ simpler or shallower networks, yield unsatisfactory results. Interestingly, despite its simplicity, WaterNet [7] performs comparatively well by leveraging preprocessed inputs (white balance, gamma correction) in addition to the original distorted image. More complex network architectures like PUGAN [14], PUIE-Net [61], and Ushape [13] achieve further performance gains. NU2Net [15], which incorporates a URanker model for underwater images into its loss function, also demonstrates strong performance. Thanks to its effectiveness, the proposed UniUIR outperforms recent state-of-the-art (SOTA) UIR methods, including Semi-UIR [26] and HCLR-Net [16]. Moreover, compared to the all-in-one SOTA method PromptIR [31] and the Mamba-based image restoration method MambaIR [43], UniUIR consistently achieves superior results across six metrics on two datasets.
To assess robustness, we also evaluated performance on non-reference datasets (Tab. III). Specifically, all methods are trained on the UIEB-T dataset and then directly tested on the non-reference datasets. In addition to the commonly used non-reference underwater image quality metrics UIQM and UCIQE, we also incorporate the URanker metric. While higher UIQM and UCIQE scores generally suggest better contrast and color vibrancy, these metrics can be biased, potentially overlooking color shifts and artifacts, as highlighted in [15]. The URanker score provides a more comprehensive evaluation of underwater image quality. Although MLLE [25], NU2Net[15], and HCLR-Net [16] tend to achieve higher UCIQE and UIQM scores, Fig. 5 reveals limitations: MLLE introduces chromatic aberrations, while NU2Net and HCLR-Net struggle with accurate color rendition, particularly noticeable in the Challenge-60 [7] and EUVP-330[55] examples. Both Semi-UIR [26] and our proposed UniUIR achieve high UIQM and UCIQE scores, along with strong URanker performance. Importantly, our method explicitly addresses the co-occurrence of multiple distortions in underwater scenes, restoring images in a unified manner. This approach enhances overall perceptual quality while preserving local details, resulting in improved contrast and vibrant colors, ultimately leading to superior performance on these challenging non-reference datasets.
TABLE IV: Effectiveness of components on the T90 [7] dataset. DAM V2 refers to Depth Anything Model V2, MMoEB to Mamba Mixture-of-Experts Block, SFPG to Spatial-Frequency Prior Generator and LCDM to Latent Condition Diffusion Model.
[…]
A (Baseline) 24.37 0.928 0.131 0.581 3.059 1.894
[…]
C 24.81 0.932 0.120 0.597 3.034 1.986
D 24.97 0.932 0.118 0.609 3.057 2.071
E 25.11 0.933 0.112 0.601 3.117 2.104
IV-C Quantitative Comparisons
This section presents a comprehensive visual comparison across reference and non-reference test sets. Beginning with the T90 dataset [7] (Fig. 5), we observe that the traditional method MLLE[25] often introduces chromatic aberrations and struggles with color distortions. Deep learning methods generally outperform MLLE, but limitations remain. For example, the WaterNet[7] struggles with yellow color casts. Ucolor[12] suffers from a loss of fine texture details. NU2Net[15] and HCLR-Net[16] occasionally introduce chromatic aberrations, as seen in the image 2 and image 4 samples of Fig. 5, indicating difficulties in controlling color. In contrast, our proposed method effectively mitigates these issues, producing enhanced images with minimal color deviations, improved contrast, and enhanced structural details. Furthermore, due to our network’s ability to address multiple distortions simultaneously, we achieve superior restoration results in challenging low-light conditions, as exemplified by image 5.
We also conduct visual comparisons across five non-reference underwater datasets. As shown in Fig. 6, many traditional and deep learning methods struggle with these datasets due to the diverse and complex distortions present. For example, MLLE frequently introduces color distortions. WaterNet and Semi-UIR struggle with haze, as seen in images 8 and 16. Additionally, WaterNet performs poorly in low-light conditions, such as image 11. NU2Net exhibits minor color casts on the Challenge-60[7], UCCS[54], and EUVP-330 datasets[55]. Ucolor and HCLR-Net struggle with yellow color casts. Furthermore, HCLR-Net performs poorly on blurry and distorted images from the U45 dataset, often lacking fine details. In contrast, our proposed UniUIR achieves the best overall visual results.
TABLE V: Effectiveness of number of top-k experts on the T90 [7] dataset.
[…]
$k$=1 24.71 0.931 0.121 0.597 3.064 1.938
$k$=2 24.94 0.932 0.117 0.602 3.025 1.995
$k$=3 25.11 0.933 0.112 0.601 3.117 2.104
$k$=4 25.01 0.932 0.115 0.607 2.993 2.052
IV-D Ablation Study
We conduct several ablation experiments to demonstrate the effectiveness of each component in the proposed UniUIR. All ablation experiments are trained on the UIEB-T [7] dataset and tested on T90 [7] dataset.
IV-D1 Effects of Different Components
As shown in Tab. IV, we evaluate the effectiveness of various components through a comparison with baseline (Index A) that excludes our modules. Integrating depth information priors derived from Depth Anything V2 demonstrably improves restoration quality, increasing PSNR, SSIM, and reducing LPIPS by 0.19 dB, 0.002, and 0.004, respectively. No-reference metrics UCIQE and URanker also show improvement. The introduction of MMoEB enhances the model’s ability to handle multiple distortions simultaneously, leading to substantial gains across both full-reference and no-reference metrics. Further incorporating SFPG and LCDM yields additional performance benefits, indicating that spatial-frequency degradation priors and latent diffusion modeling contribute to more effective underwater image restoration. Notably, UCIQE and UIQM trends don’t always align with other metrics (e.g., UIQM in Indexs A and B, and UCIQE in Indexs D and E), potentially echoing findings in [15] regarding limitations of UCIQE and UIQM in reflecting perceptual underwater image quality. Finally, the visual results in Fig. 7 demonstrate that our proposed approach (Index E) achieves the best visual quality.
IV-D2 Effects of Number of Top-k Experts
Tab. V presents the effectiveness of varying the number of top-$k$ experts on the T90 dataset. It can be observed that the trends for PSNR, SSIM, LPIPS, and URanker are consistent, all achieving their best values when $k=3$. Although the trends for UCIQE and UIQM differ slightly from the aforementioned metrics, the best UIQM value is also obtained at $k=3$. These findings suggest that selecting $k=3$ experts strikes the optimal balance across all metrics, making it the most suitable configuration for the experiments.
TABLE VI: Performance comparisons on low-light and backlit image enhancement tasks.
Method LOL-v1 [63] LOL-v2-real [64] SID [65] BAID [66]
[…]
URetinex [67] 21.33 0.835 85.59 20.44 0.806 76.74 22.09 0.633 71.58 19.08 0.845 42.26
Restormer [42] 22.43 0.823 78.75 19.94 0.827 114.35 22.27 0.649 75.47 21.07 0.832 41.17
SNR-Net [68] 24.61 0.842 66.47 21.48 0.849 68.56 22.87 0.625 74.78 20.86 0.860 39.73
DiffIR [45] 23.15 0.828 70.13 21.15 0.816 72.33 23.17 0.640 78.80 21.10 0.835 40.35
Diff-Retinex [40] 21.98 0.852 51.33 20.17 0.826 46.67 23.62 0.665 58.93 22.07 0.861 38.07
UniUIR 24.83 0.851 55.19 22.69 0.850 46.23 24.67 0.685 60.64 22.21 0.864 36.23
Figure 8: Qualitative results on LOL-v1 (top) , LOL-v2-real (middle), and BAID (bottom).
Figure 9: The depth maps and semantic segmentation results generated by Depth Anything V2 [21] and the Segment Anything Model [69] are presented in Rows 2 and 4, respectively. These images are derived from the original underwater images input and those restored using Ucolor, NU2Net, GUPDM, HCLR-Net, the proposed UniUIR, and corresponding reference images. For ease of comparison, the restored images are also displayed in Rows 1 and 3.
Figure 10: Demonstration of failure cases for the proposed UniUIR and state-of-the-art underwater image restoration methods on the T90 dataset.
IV-E Generalization
IV-E1 Application on Other Low-level Vision Tasks
To validate the effectiveness of the proposed UniUIR for low-light and backlit image enhancement tasks, we retrained UniUIR and compared its performance against state-of-the-art image enhancement methods on the LOL-v1 [63], LOL-v2-real [64], SID [65], and BAID [66] datasets. As shown in Tab. VI, UniUIR achieves state-of-the-art performance across all datasets, outperforming existing methods in terms of PSNR, SSIM, and perceptual quality metrics. This demonstrates the robustness and generalizability of UniUIR in handling complex low-light and backlit scenarios. Fig. 8 provides qualitative results, highlighting our ability to produce enhanced images with corrected illumination and enhanced texture in low-light and backlit scenes. Existing enhancement methods often struggle in scenarios involving intricate lighting conditions or high-frequency textures. UniUIR addresses these challenges through its adaptive feature extraction and spatial-frequency prior generation, ensuring more accurate illumination correction and texture enhancement. These capabilities enable UniUIR to deliver consistently high-quality results, even in challenging regions.
IV-E2 Application on High-level Vision Tasks
We investigate the impact of restored underwater images on downstream tasks, including depth estimation and image segmentation. i) For depth estimation task, we employed the Depth Anything V2 [21] model to extract depth maps from restored images and compared the performance of various UIR methods. As illustrated in Fig. 9 (orange and red boxes in the second row), UniUIR produced depth maps that closely matched those suggested by the reference image, outperforming previous methods in handling depth-related distortions. This indicates that UniUIR better preserves depth consistency and enhances performance on depth estimation tasks. ii) For image segmentation task, we used enhanced images from the T90 dataset produced by various UIR methods and directly applied the Segment Anything Model (SAM) [69] to evaluate the impact of underwater image restoration on segmentation performance. Notably, as shown in Fig. 9 (fourth row, yellow dashed box), only HCLR-Net [16] and our UniUIR correctly identified the background diver’s fins, despite occlusion by foreground fish. However, the visual quality of HCLR-Net’s restoration is inferior to that of UniUIR. Interestingly, in the red dashed box (fourth row, Fig. 9), the reference image’s segmentation erroneously merged two distinct fish into one, while all UIR methods, including the distorted input, successfully distinguished them. This highlights the importance of balancing human-perceived quality and machine vision performance for downstream tasks.
These results demonstrate that UniUIR not only achieves superior visual quality but also significantly improves the utility of restored underwater images for machine vision tasks. The findings underscore the necessity of designing UIR methods that optimize both human-perceived quality and downstream task performance.
IV-F Limitation and Future Work
Although our proposed method consistently outperforms state-of-the-art approaches in underwater image enhancement, certain limitations persist. As illustrated in Fig. 10, existing methods struggle with enhancing underwater images under extreme low-light and blurry conditions. This difficulty can be attributed to the fact that most current datasets emphasize common distortions such as color shifts and low contrast, with limited examples of extreme scenarios like those encountered in low-light environments. Consequently, the learning capacity of models is constrained, leading to challenges in handling more complex conditions. Moreover, our method leverages the pre-trained Depth Anything V2 model, which has been primarily trained on natural scenes containing minimal underwater data. Directly applying its pre-trained weights may result in suboptimal capture of depth features specific to underwater settings.
To address these issues, future work could employ image generation techniques to augment the training dataset with synthetic extreme underwater scenes. This would enrich the training set with diverse and realistic samples, significantly enhancing the model’s generalization ability in complex scenarios. Furthermore, exploring efficient fine-tuning strategies for depth extraction tailored to underwater environments can be highly beneficial. For example, employing Low-Rank Adaptation (LoRA) techniques to fine-tune pre-trained depth extraction models [70] would enable the recovery network to better accommodate underwater-specific characteristics such as light variations, turbidity, and depth inconsistencies. This approach would enhance the model’s robustness and performance under challenging underwater conditions.
V Conclusion
In this paper, we systematically review the current challenges in underwater image restoration and propose UniUIR, an all-in-one approach designed to address the complex underwater scenario. To effectively harness the interdependencies among various degradations in underwater environments, we developed the Mamba Mixture-of-Experts module. This module allows each expert to specialize in recognizing specific types of degradation and collaboratively infer task-specific priors while preserving both local and global feature representations. To enhance the model’s ability to identify inconsistencies between distortions across different image regions, we integrate scene depth information using the Depth Anything V2 model. Additionally, we propose a spatial-frequency prior generator that extracts degradation priors from both spatial and frequency domains via task-adaptive routing. This mechanism enables conflicting tasks to utilize distinct network paths, thereby reducing task interference. Compared to state-of-the-art underwater image restoration methods, UniUIR delivers superior results, as evidenced by extensive qualitative and quantitative evaluations.